# Out of spuriousity: Improving robustness to spurious correlations without group annotations

### v1

@@ 5.3 Overall Training Procedure
Given a trained ERM neural network $f$ on dataset $\mathcal{D}$, we perform the following steps sequentially.
Representation clustering and sub-sampling. First, we extract the representations learned by the ERM model as the output of the last hidden layer of $f$: $f_{emb}$. We apply $k-$means on $f_{emb}(x)$ with a pre-defined $k$. We determine cluster labels ($i-$dominant and neutral) and construct the subnetwork mask-training dataset $\mathcal{D}_{task}$ as described in Sec. 5.2.2. In our experiments, we set $p$, the number of samples for each class such that $\mathcal{D}_{task}$ typically contains 10% of the samples of $\mathcal{D}$.
Binary mask training. We freeze all weights $\mathbf{W}$ of $f$ and train only the mask $\bm{\Pi}$ with the mask-training dataset $\mathcal{D}_{task}$ (cf. Sec. 4). To construct contrastive batches, we randomly sample multiple anchors $x$ from distinct cluster together with their corresponding positive $x^{+}$ and negative $x^{-}$. We accumulate the per-batch constrastive loss by summing over $\mathrm{L_{con}}$ of all anchors in the batch. The mask $\bm{\Pi=(\pi^{1},\pi^{2},...,\pi^{L})}$ is updated through the final loss term
$\mathrm{L}=\mathrm{L_{CE}}+\alpha\sum_{i}\pi_{i}+\beta\mathrm{L_{con}},$
@@
After binarization of the trainable mask $\bm{\Pi}$, the resulting binary mask $\mathbf{m}$ defines the parameters of the subnetwork as $\mathbf{W^{\prime}=(w^{1}\odot m^{1},w^{2}\odot m^{2},...,w^{L}\odot m^{L})}$. We finetune the extracted subnetwork on the dataset $\mathcal{D}_{task}$ with cross-entropy loss for a few epochs.
@@ 6 Experiments
We evaluate the effectiveness of PruSC on three datasets (CelebA, WaterBird, and ISIC) in comparison to state-of-the-art methods (Sec. 6.3). We further examine the impact of PruSC in the presence of multiple potential spurious correlations (Sec. 6.4) and conduct an ablation study for the main components of our approach (Sec. 6.5).
@@ 6.1 Datasets
Following prior work, our classification target is hair color (blond or non-blond hair) on the CelebA dataset (Liu et al., 2015), where the potential spurious attribute is gender (male or female). Majority groups are blond females and non-blond males with proportions of $44\%$ and $41\%$ of the data. Minority groups are non-blond females ($14\%$) and blond males ($1\%$). We extend the common setup to a multi-spuriouss dataset by incorporating additional attributes (no beard, young, heavy makeup, wearing lipstick, pale skin) which also pose potential spurious correlations based on their heavily imbalanced distributions.
WaterBird.
WaterBird was introduced by Sagawa et al. (2020a) as a standard spurious correlation benchmark. This artificially constructed dataset pastes bird segmentations from the CUB dataset (Wah et al., 2011) onto backgrounds form the Places dataset (Zhou et al., 2018). The birds labeled as either waterbirds or landbirds are placed on backgrounds of either water or land. The task is to classify bird types: $\mathcal{Y}$ = {waterbird, landbird} and pasting $95\%$ waterbirds on a water background and $95\%$ landbirds on a land background results in spurious correlations in the default training set since models tend to rely on the background instead of actual bird characteristics. Accordingly, the spurious attributes are $\mathcal{A}$ = {water background, land background}.
@@
Additional details about the distribution of datasets with respect to spurious features are provided in Appendix A.1.
@@ 6.2 Baselines
Empirical Risk Minimization (ERM) represents conventional training without any procedures for improving worst-group accuracies.
We compare PruSC with two approaches that require group annotations: GroupDRO (Sagawa et al., 2020a) needs knowledge about spurious attributes and group annotations to train from scratch, while DFR (Kirichenko et al., 2023) requires a group-balanced subset for its retraining stage. To meet this requirement in the ISIC dataset, we create artificial samples for the missing group and use those samples during the training of GroupDRO and DFR. To highlight the need of group annotations in DFR, we implement DFRclass, which is DFR but uses our automatically inferred class-balanced subset $\mathcal{D}_{task}$ from clustering instead of annotated group-balanced data.
Additionally, we compare to methods with relaxed annotation requirements. CnC (Zhang et al., 2022) and DCWP (Park et al., 2023) work with ERM misclassified cases instead of group annotated samples. These methods sometimes necessitate retraining ERM with early stopping criteria to obtain a sufficiently large set of misclassified samples. DeDiER (Tiwari et al., 2024), is an early readout mechanism with a distillation model. This method identifies misclassified cases from early layers of the network by design. However, DeDiER requires group information during validation to tune hyper-parameters and determine the early layer from which to take the readout.
In contrast to state-of-the-art methods in our comparison, PruSC does not require group annotations at all, neither for (re-)training, nor for fine-tuning hyper-parameters (cf. annotation requirements in the first column of Tab. 2).
DCWP, DFR, DFRclass and PruSC, first require an ERM model influenced by spurious correlations before taking actions to mitigate those spurious correlations. For fair comparison, we use the same ERM model which is the ERM baseline reported in Tab. 2 for all methods. Since the training accuracy of this baseline ERM is around $99\%$, DCWP requires another early stopped model for gathering a sufficient amount of misclassified cases.
@@
We use ResNet18 and ResNet50 models pretrained on ImageNet. We train the models for 50 epochs with SGD with a constant learning rate of $10^{-3}$, momentum decay of 0.9, weight decay of $10^{-2}$, and batch size $32-64$. All images are resized to $224\times 224$. We use ResNet50 for the predictive performance evaluation in Tab. 2 to be comparable to the original work for CnC and DeDiER. We use ResNet18 to investigate the robustness of our model to multiple spurious correlations (Sec. 6.4) and in the ablation study (Sec. 6.5). For subnetwork extraction approaches (DCWP and PruSC), we report results with pruning around 50% of the network parameters and discuss the impact of other pruning ratios in Sec. 6.7.
@@ 6.3 Predictive Performance
Table 2: Performance overview, comparing methods with different assumptions on availability of annotations (A.) of the spurious features in training (Tr) and/or validation set (Val). ∗∗ denotes results reported in respective papers, n.a. – results not available in the original paper. denotes the presence of an artificially created missing group in ISIC, as otherwise DFR and GroupDRO would not be applicable. We bold the highest WGA without any group annotation requirement and underline the best among all methods. We report the mean and standard deviation over 5 runs.
w/o A. CelebA ISIC WaterBird
@@
✓/✓ ERM (baseline) 90.1$\pm$0.3 49.7$\pm$1.0 86.9$\pm$0.2 34.4$\pm$0.8 89.5$\pm$0.2 63.3$\pm$0.4
✗/✗ DFR 91.3$\pm$0.3 88.3$\pm$1.1 87.4$\pm$1.3 77.7$\pm$2.4 95.6$\pm$0.5 85.6$\pm$0.4
✗/✗ GroupDRO 93.0$\pm$0.1 88.7$\pm$0.3 87.7$\pm$0.8 59.0$\pm$0.4 90.3$\pm$0.0 87.0$\pm$0.1
✓/✗ CnC∗∗ 93.9$\pm$0.1 88.9$\pm$1.3 n.a. n.a. 92.0$\pm$0.6 89.9$\pm$0.6
✓/✗ DeDiER∗∗ 93.2$\pm$0.06 89.6$\pm$1.67 n.a n.a 92.1$\pm$0.39 89.8$\pm$0.47
✓/✓ DCWP 88.9$\pm$2.3 73.2$\pm$2.0 86.4$\pm$1.7 45.1$\pm$0.8 85.0$\pm$0.9 54.8$\pm$2.1
✓/✓ DFRclass 84.9$\pm$2.1 67.7$\pm$0.8 74.1$\pm$2.6 48.8$\pm$6.0 87.5$\pm$1.2 74.6$\pm$2.9
✓/✓ PruSC 91.0$\pm$0.3 89.7$\pm$0.3 86.1$\pm$0.4 75.1$\pm$1.7 90.3$\pm$2.2 79.8$\pm$0.8
Our main results on predictive performance are shown in Tab. 2. PruSC shows the best performance in terms of worst-group accuracy (WGA) among methods that do not require any group annotations and shows comparable performance in most cases (CelebA and ISIC) to state-of-the-art methods that rely on annotations. Notably, our method successfully improves the accuracy of the worst group malignant with patch in ISIC, although it has never seen an example of that group during training. Our performance is even comparable with GroupDRO and DFR which explicitly need (artificial) minority group samples during training.
The DFRclass method mimics DFR but replaces the group-balanced subset (which requires manual annotations) by a class-balanced subset (class annotations are readily available). DFRclass falls short compared to DFR in all cases by a large margin, highlighting the importance of group annotations in DFR retraining. Using the same retraining dataset as DFRclass, the success of our method PruSC emphasizes the necessity of distorting spurious clusters, which we achieve through a contrastive loss. We empirically analyze the impact of indiviual components of our method in an ablation study (Sec. 6.5).
While our approach has comparable performance to state-of-the-art on the CelebA and ISIC dataset, it seems to be less effective on the WaterBird dataset. We identified two potential factors, that could explain this behavior. First, the WaterBird dataset has spurious-free validation and test sets, where each class has equal quantities of land and water backgrounds. Therefore, methods that utilize the validation set during training can improve performance without needing explicit group labels. Second, the WaterBird dataset is constructed from 17 bird species divided into two categories: waterbirds and landbirds. Our approach samples a subset $\mathcal{D}_{task}$ of only around 10% of the training set without access to the bird species information. Consequently, we might lose crucial bird information during sub-sampling, and mask weights containing information that, while irrelevant to the functional task defined on $\mathcal{D}_{task}$, is essential for recognizing bird types.
@@ 6.4 Robustness to Multiple Spurious Correlations
In this section, we analyze how PruSC performs in the presence of multiple spurious correlations within a dataset. Since we do not explicitly use information about the spurious attribute during subnetwork training, we expect the representations to be agnostic to any spurious feature. We use the CelebA dataset which has a potential set of spurious attributes $\mathcal{A}$ (Seo et al., 2022) (for details see Appendix A.3). Specifically, we investigate six potential spurious attributes $a_{i}\in\mathcal{A}=$male, presence of beard, heavy makeup, wearing lipstick, young, pale skin}. For each attribute $a_{i}$, we create a group set $\mathcal{G}$ = {(blond, $a_{i}$), (non-blond, $a_{i}$), (blond, non-$a_{i}$), (non-blond, non-$a_{i}$)}. We train an ERM model to predict hair color and evaluate performance separately for each group in $\mathcal{G}$. We then train all methods for correcting spurious attributes. GroupDRO and DFR are trained on balanced groups for the attribute male. [^3] GroupDRO and DFR are single-attribute methods, and cannot be trained for multiple attributes at the same time. Since the spurious attributes in CelebA are themselves strongly correlated (male, beard) there are still performance improvements for attribute groups not specifically trained for.
6.4.1 Evaluation Metrics
We report worst-group accuracy (WGA) for each spurious attribute $a_{i}$ evaluated on groups in $\mathcal{G}$. WGA is the standard metric to evaluate whether models use spurious correlations. A significant drop between average accuracy and worst-group accuracy is a strong indicator of relying on spurious correlations. To assess the impact on the group with the fewest training samples (usually the group with a spurious attribute uncorrelated to a target label), we report minority-group accuracy (MGA) for each spurious attribute $a_{i}$. To assess group fairness, i.e., whether all groups have similar prediction error, we measure the unbiased accuracy gap (UAG). Unbiased accuracy (UA) is the accuracy on a balanced test set, sampled such that there is no correlation between the attribute $a_{i}$ and a target label. UGA is the difference of the overall (official) test set accuracy and UA. A small UAG indicates that groups are treated equally, regardless of their distribution in the dataset.
6.4.2 Results
Table 3: Minority group accuracy across multiple spurious attributes in CelebA. Each method train with CelebA with target $\mathcal{T}=$blond hair}, spurious attributes set are $\mathcal{A}=$male (M), no beard (NB), heavy makeup (HM), wearing lipstick (WL), young (Y), pale skin (PS)}. GroupDRO and DFR train extensively for attribute $a_{i}=$ M while ERM (baseline), DCWP, and Ours only train with target label information. We report the performances of the minority groups defined in the training set - those groups are aligned with the worst-group defined by ERM worst-group accuracy. We bold the largest increase from the baseline among all methods.
w/o A. M NB HM WL Y PS
@@ 6.4.2 Results
✓/✓ ERM (baseline) 50.2 37.5 79.6 79.4 87.5 83.5
✗/✗ GroupDRO 88.8 85.9 90.6 89.9 91.3 87.5
✗/✗ DFR 85.6 78.1 83.1 82.6 85.4 81.9
✓/✓ DCWP 81.8 78.1 76.9 76.9 83.7 79.2
✓/✓ PruSC 91.2 92.2 92.7 92.2 92.8 88.7
The ERM models usually fail for the minority group represented by target and uncorrelated attributes. For example, for the attribute gender = male, the ERM model only has $50.2\%$ accuracy for the group (blond hair, male) since in the training set there are only $1,102$ samples in the group (blond hair, male) while the group (non-blond hair, male) contains $53,483$ samples. In all of these cases, the accuracy for the minority group is the worst-group accuracy evaluated by the baseline ERM model.
In Tab. 3, we report the performance of the minority group across multiple spurious attributes. The details of data distribution are shown in Tab. 7 and the full details for group accuracy of the groups $\mathcal{G}$ = {(blond, $a_{i}$), (non-blond, $a_{i}$), (blond, non-$a_{i}$), (non-blond, non-$a_{i}$)} in Tab. 9) (Appendix).
Tab. 3 shows that PruSC has the largest improvement from the baseline in all spurious attribute cases. Especially in the case of the spurious feature “male”, PruSC performs better than GroupDRO or DFR which are trained with known group labels for this attribute.
The worst-group accuracy of baselines ERM, using group labels GroupDRO and PruSC not using prior knowledge of spurious attributes are reported in Tab. 4 (WGA).
The increase in worst-group accuracy (WGA) compared to the baselines suggests that all methods are able to unlearn certain biased features. However, in a realistic dataset, e.g. CelebA, there are identical distributions in the train and official test set a model can use to obtain high accuracy on the majority group, but fail on other groups. Therefore, measuring WGA on the official test set can prove that the methods decrease the impact of bias features, but can not tell to what extent the model treats all groups equally. Tab. 4 shows that our approach is on par with GroupDRO in UA and WGA. This indicates that both methods can effectively unlearn spurious correlations and maintain strong performance even when there is a distribution shift from a highly imbalanced training set to a balanced test set. However, PruSC shows a smaller UGA in all cases and is, therefore, more consistent among different spurious attributes and shows a more balanced performance among groups.
Table 4: Multi-attributes. The table reports unbiased accuracy (UA), worst-group accuracy (WGA) and unbiased accuracy gap (Gap) - the gap between average and corresponding unbiased accuracy. We bold the highest WGA and underline the smallest Gap for each spurious attribute. The $\uparrow$ indicates the higher the better and the $\downarrow$ indicates the lower the better.
ERM GroupDRO PruSC
@@
male 81.6 50.2 7.8 90.7 88.5 2.2 90.5 89.6 0.4
no beard 79.8 37.5 9.6 90.9 85.9 2.0 91.0 89.6 0.9
heavy makeup 90.1 79.6 0.7 91.4 89.4 1.5 90.2 88.9 0.1
wearing lipstick 88.8 79.4 0.6 90.9 89.6 2.0 90.1 88.7 0.0
young 91.8 87.5 2.4 91.0 84.5 1.9 89.3 81.4 0.8
pale skin 91.4 83.5 2.0 90.8 87.5 2.1 89.8 88.7 0.3
6.5 Ablation study
Tab. 5 shows an overview of our ablations: performing the pruning and fine-tuning steps, with and without our contrastive loss, respectively (ConLoss Pr, ConLoss Ft).
Contrastive Loss.
Settings 1 and 2 demonstrate the impact the contrastive loss during fine-tuning (ConLoss Ft) when training a subnetwork. Similarly, settings 5 and 6 show the influence of contrastive loss while not pruning but only fine-tuning the whole network with $\mathcal{D}_{task}$. In both cases, using the contrastive loss in fine-tuning downgrades the performance. Setting 3 reveals that without contrastive loss in the training mask (ConLoss Pr), the WGA drops significantly (89.6% from setting 1 to 41.8% in setting 3). It shows that our final combination (only using contrastive loss during pruning) is the best combination and surpasses other combinations by large margins. In other words, our defined contrastive loss is most effective when applied to the mask training instead of directly to the model weight training. The contrastive loss forces the pruning process to retain only the connections that contribute to an optimized representation, ensuring that all spurious attributes are distributed in the representation space. Furthermore, in the pruning stage, the backpropagation process only updates the masks, which later will be binarized by Gumbel-Sigmoid, instead of directly updating the model weights.

### v2

@@ 5.3 Overall Training Procedure
Given a trained ERM neural network $f$ on the dataset $\mathcal{D}$, we perform three steps sequentially.
Representation Clustering and Sub-sampling.
First, we extract the representations learned by the ERM model as the output of the last hidden layer of $f$: $f_{emb}$. We apply $k$-means to $f_{emb}(x)$ with a pre-defined kk. [^4] We use $k=8$ for all experiments and show that the number of clusters not a critical hyper-parameter for our approach, as long as it is large enough in section 7.2 We determine cluster labels (ii-dominant and neutral) and construct the subnetwork mask-training dataset $\mathcal{D}_{\text{task}}$ as described in Sec. 5.2.2. In our experiments, we set $p$, the number of samples for each class such that $\mathcal{D}_{\text{task}}$ typically contains 10% of the samples of $\mathcal{D}$.
Binary Mask Training.
We freeze all weights $\mathbf{W}$ of $f$ and train only the mask $\bm{\Pi}$ with the mask-training dataset $\mathcal{D}_{\text{task}}$ (cf. Sec. 4). To construct contrastive batches, we randomly sample multiple anchors $x$ from different clusters along with their corresponding positives $x^{+}$ and negatives $x^{-}$. We accumulate the per-batch constrastive loss by summing over $\mathrm{L_{con}}$ of all anchors in the batch. The mask $\bm{\Pi=(\pi^{1},\pi^{2},...,\pi^{L})}$ is updated through the final loss term
$\mathrm{L}=\mathrm{L_{CE}}+\alpha\sum_{i}\pi_{i}+\beta\mathrm{L_{con}},$ (7)
@@
After binarization of the trainable mask $\bm{\Pi}$, the resulting binary mask $\mathbf{m}$ defines the parameters of the subnetwork as $\mathbf{W^{\prime}=(w^{1}\odot m^{1},w^{2}\odot m^{2},...,w^{L}\odot m^{L})}$. We fine-tune the extracted subnetwork on the dataset $\mathcal{D}_{\text{task}}$ with cross-entropy loss for a few epochs.
@@ 6 Experiments
We evaluate the effectiveness of PruSC on three datasets (CelebA, Waterbirds, and ISIC) in comparison to state-of-the-art methods (Sec. 6.4). We further examine the impact of PruSC in the presence of multiple potential spurious correlations (Sec. 6.5).
@@ 6.1 Datasets
Following previous work, our classification target is hair color (blond or non-blond hair) on the CelebA dataset (Liu et al., 2015), where the potential spurious attribute is gender (male or female). Majority groups are blond females and non-blond males with proportions of $44\%$ and $41\%$ of the data. The minority groups are non-blond females ($14\%$) and blond males ($1\%$). We extend the common setup to a multi-spurious dataset by incorporating additional attributes (no beard, young, heavy makeup, wearing lipstick, pale skin) that also pose potential spurious correlations based on their heavily imbalanced distributions.
Waterbirds.
Waterbirds was introduced by Sagawa et al. (2020a) as a standard spurious correlation benchmark. This artificially constructed dataset pastes bird segmentations from the CUB dataset (Wah et al., 2011) onto backgrounds from the Places dataset (Zhou et al., 2018). The birds labeled as either waterbirds or landbirds are placed on backgrounds of either water or land. The task is to classify the types of birds: $\mathcal{Y}$ = {waterbird, landbird} and pasting $95\%$ waterbirds on a water background and $95\%$ landbirds on a land background results in spurious correlations in the standard training set since the models tend to rely on the background instead of the actual bird characteristics. Accordingly, the spurious attributes are $\mathcal{A}$ = {water background, land background}.
@@
Details about the distributions of spurious attributes in each dataset are provided in Appendix A.1.
@@ 6.2 Baselines
Empirical Risk Minimization (ERM) represents conventional training without any procedures to improve worst-group accuracy.
We compare PruSC with two approaches that require group annotations: GroupDRO (Sagawa et al., 2020a) requires knowledge about spurious attributes and group annotations to train from scratch, while DFR (Kirichenko et al., 2023) requires a group-balanced subset for its retraining phase. To meet this requirement in the ISIC dataset, we create artificial samples for the missing group (malignant with patches) according to Nauta et al. (2021). We use these artificial samples during the training of GroupDRO and DFR. To analyze the importance of a group-balanced subset in DFR, we implement DFRclass, which is a version of DFR that uses our automatically inferred class-balanced subset $\mathcal{D}_{task}$ from clustering instead of annotated group-balanced data.
In addition, we compare to methods with relaxed annotation requirements. CnC (Zhang et al., 2022) and DCWP (Park et al., 2023) work with ERM misclassified cases instead of group annotated samples. These methods sometimes necessitate retraining the ERM with early stopping criteria to obtain a sufficiently large set of misclassified samples. SPARE (Yang et al., 2024) shows that groups in $\mathcal{G}$ can be separated early in ERM training if the spurious correlations are strong enough. It infers groups from early training epochs and from these inferred groups samples group-balanced mini batches during training of the remaining epochs. DeDiER (Tiwari et al., 2024), is an early readout mechanism with a distillation model. This method identifies misclassified cases from early layers of the network by design. However, both SPARE and DeDiER require group information during validation to tune the hyper-parameters and determine the early epoch or layer from which to take the readout. Early-stop disagreement SELF (LaBonte et al., 2024) fine-tunes the last layer weights with a small held-out dataset chosen by maximizing the disagreement between ERM model and an early-stop ERM model. SELF requires neither group annotations nor class labels for the held-out dataset, however, it does require group annotation during hyper-parameters tuning. In contrast to the above methods, PruSC does not require any group annotations, neither for (re-)training, nor for fine-tuning hyper-parameters (cf. annotation requirements in the first column of Tab. 2).
DCWP, DFR, DFRclass and PruSC, first require an ERM model that is affected by spurious correlations before taking action to mitigate those spurious correlations. For a fair comparison, we use the same ERM model, which is the ERM baseline reported in Tab. 2 for all methods. Since the training accuracy of this baseline ERM is about $99\%$, DCWP needs access to the full training process of the ERM model and uses earlier model checkpoints to obtain a sufficient amount of misclassified samples.
@@
We use ResNet18 and ResNet50 pre-trained on ImageNet. We train the models for 50 epochs using SGD with a constant learning rate of $10^{-3}$, momentum decay of 0.9, weight decay of $10^{-2}$, and batch size 326432-64. All images are resized to $224\times 224$. We use ResNet50 to evaluate the predictive performance in Tab. 2 to compare with the original work for CnC, DeDiER and SELF. We use ResNet18 to investigate the robustness of our model to multiple spurious correlations (Sec. 6.5) and in the ablation study (Sec. 7.1). For the subnetwork extraction approaches (DCWP and PruSC), we report results with pruning around 50% of the network parameters and discuss the impact of other pruning ratios in Sec. 7.3.
6.3 Evaluation Metrics
We report worst-group accuracy (WGA), the accuracy of the group (attribute-label combination) that a particular method performs worst. WGA is the standard metric for evaluating whether models use spurious correlations (Sagawa et al., 2020a). A significant drop between average accuracy and worst-group accuracy is a strong indicator of reliance on spurious correlations. To assess the impact on the group with the fewest training samples (usually the group with a spurious attribute uncorrelated with a target label), we report minority-group accuracy (MGA) for each spurious attribute $a_{i}$. The difference between WGA and MGA is that the former is defined per method, and the latter is defined on the data. Since the minority group is usually the group with a spurious attribute uncorrelated with a target label, performance is mostly lowest on this group and WGA==MGA, but not necessarily. To assess group fairness, i.e., whether all groups have similar prediction errors, we measure the unbiased accuracy gap (UAG). Unbiased accuracy (UA) is the accuracy on a balanced test set, sampled such that there is no correlation between the attribute $a_{i}$ and a target label. UAG is the difference between the overall (official) test set accuracy and the UA. A small UAG indicates that groups are treated equally, regardless of whether they are minority or majority groups.
@@ 6.4 Predictive Performance
Our main results on predictive performance are shown in Tab. 2. PruSC shows the best performance in terms of worst-group accuracy (WGA) among methods that do not require any group annotations and shows comparable performance in most cases (CelebA and ISIC) to state-of-the-art methods that rely on annotations. Notably, our method successfully improves the accuracy of the worst group malignant with patch in ISIC, although it has never seen an example of this group during training. Our performance is even comparable to GroupDRO and DFR which explicitly need (artificial) minority group samples during training.
Table 2: Performance overview, comparing methods with different assumptions on availability of annotations (Annot.) of the spurious features in training (Tr) and/or validation set (Val). ∗∗ denotes results reported in respective papers, n.a. denotes that results are not available in the original paper. denotes the presence of an artificially created missing group in ISIC, as otherwise DFR and GroupDRO would not be applicable. We bold the highest WGA without any group annotation requirement and underline the best among all methods. We report the mean and standard deviation over 5 runs.
Annot. CelebA ISIC Waterbirds
@@
[…]
The DFRclass method mimics DFR but replaces the group-balanced subset (which requires manual annotations) with a class-balanced subset (class annotations are readily available). DFRclass falls short compared to DFR in all cases by a large margin, highlighting the importance of group annotations in DFR retraining. Using the same retraining dataset as DFRclass, the success of our method PruSC emphasizes the necessity of distorting spurious clusters, which we achieve through a contrastive loss. We empirically analyze the impact of individual components of our method in an ablation study (Sec. 7.1).
On Waterbirds, there is a clear gap between annotation-free methods and those that rely on annotations. Using only annotations in the validation set, SELF even outperforms methods that require more annotations. The reason is a peculiarity of the Waterbirds dataset: its validation set is spurious-free, i.e., each class has equal quantities of land and water backgrounds, which gives an advantage to methods that rely on the validation set. This advantage becomes particularly apparent for DFRclass. By sampling a class-balanced subset from the validation set instead of the training set, LaBonte et al. (2024) report a WGA of $92.6$ for DFRclass, which is comparable to the performance of DFR on a group-balanced subset obtained from label annotations. For truly annotation-free methods (such as ours) a spurious-free validation is not available, as its construction itself requires annotations.
@@ 6.5 Robustness to Multiple Spurious Correlations
In this section, we analyze how PruSC performs in the presence of multiple spurious correlations within a dataset. Since we do not explicitly use information about the spurious attribute during subnetwork training, we expect the representations to be agnostic to any spurious feature. We use the CelebA dataset which has a potential set of spurious attributes $\mathcal{A}$ (Seo et al., 2022) (see Appendix A.3 for details). Specifically, we investigate six potential spurious attributes $a_{i}\in\mathcal{A}=$ {male, presence of beard, heavy makeup, wearing lipstick, young, pale skin}. For each attribute $a_{i}$, we create a group set $\mathcal{G}$ = {(blond, $a_{i}$), (non-blond, $a_{i}$), (blond, non-$a_{i}$), (non-blond, non-$a_{i}$)}. We train an ERM model to predict hair color and evaluate the performance separately for each group in $\mathcal{G}$. We then train all methods for correcting spurious attributes. GroupDRO and DFR are trained on balanced groups for the attribute male. [^7] GroupDRO and DFR are single-attribute methods, and cannot be trained for multiple attributes at the same time. Since the spurious attributes in CelebA are themselves strongly correlated (male, beard) there are still performance improvements for attribute groups not specifically trained for.
In Tab. 3, we report the performance of the minority group across multiple spurious attributes. Details of the data distribution and accuracy values for each individual group $\mathcal{G}$ = {(blond, $a_{i}$), (non-blond, $a_{i}$), (blond, non-$a_{i}$), (non-blond, non-$a_{i}$)} are provided in the Appendix, Tab. 7 and Tab. 9.
Table 3: Minority group accuracy (MGA) across multiple spurious attributes $\mathcal{A}=$ {male, no beard, heavy makeup, wearing lipstick, young, pale skin} on CelebA with classification target hair color. ERM (baseline), DCWP and PruSC are trained with class labels only, GroupDRO and DFR in addition on attribute $a_{i}=$ male to mitigate the spurious correlation on gender. We report the accuracy of the minority groups defined in the training dataset. We bold the highest MGA among all methods.
Annot. male no beard heavy makeup wearing lipstick young pale skin
@@
[…]
The ERM models usually fail for the minority group represented by the target and uncorrelated attributes. For example, for the attribute gender = male, the ERM model only has $48.5\%$ accuracy for the group (blond hair, male) since in the training set there are only $1,102$ samples in the group (blond hair, male) while the group (non-blond hair, male) contains $53,483$ samples. For the ERM model in Tab. 3, the minority group accuracy is indeed always equal to worst group accuracy, which is not the case for the other methods. For example, GroupDRO has an accuracy of $88.7\%$ on the minority group (blond hair, male), but worst-group accuracy of $86.7\%$ for the group (not blond hair, not male). PruSC has the largest improvement over the baseline across all spurious attributes. In particular, on the spurious attribute “male”, PruSC even outperforms GroupDRO and DFR which were trained with known group labels for this attribute.
Table 4: Multiple spurious attributes. We report unbiased accuracy (UA), worst-group accuracy (WGA) and unbiased accuracy gap (UAG). We bold the highest WGA and underline the smallest UAG for each spurious attribute. $\uparrow$ higher is better, $\downarrow$ lower is better.
male no beard heavy makeup wearing lipstick young pale skin
[…]
In Tab.4 we compare GroupDRO that relies on group annotations, with PruSC that does not require prior knowledge of spurious attributes. Both methods show an increase of worst-group accuracy (WGA) over the ERM baseline for all groups, suggesting that both methods can successfully mitigate certain spurious correlations. However, in CelebA’s official training and test split, the groups are identically distributed. The WGA indicates whether a method has successfully reduced the impact of biasing features, but because the average accuracy is driven by the majority group, it may mask performance declines in the remaining groups. That is, we cannot determine from these two measures alone whether a model treats all groups equally. Our approach, PruSC, is on par with GroupDRO in terms of unbiased accuracy (UA) and WGA. This indicates that both methods can effectively unlearn spurious correlations and maintain strong performance even when there is a distribution shift from a highly imbalanced training set to a balanced test set. However, PruSC shows a smaller UGA in all cases and is therefore more consistent across different spurious attributes and shows a more balanced performance across groups.
7 Detailed Analysis
In this section, we show an ablation study to evaluate the effectiveness of PruSC components, including the pruning and fine-tuning steps, as well as the use of contrastive loss (Sec.7.1). We also analyze the impact of number of clusters (Sec.7.2), the impact of varying pruning ratios (Sec.7.3), the performance with different choices of contrastive batch in PruSC (Sec.7.4) and the representation space after mitigation (Sec.7.5).
7.1 Component Ablation
We analyze the impact of each component of PruSC. Tab. 5 shows an overview of our settings. Specifically, we evaluate the impact of pruning, i.e., subnetwork extraction and fine-tuning on performance, along with the effect of applying contrastive loss at each step. If the contrastive loss is omitted during pruning (ConLoss Pr ✗), we train the subnetwork with cross-entropy loss and sparse regularization. Without the contrastive loss during fine-tuning (ConLoss Ft ✗), we fine-tune the subnetwork or full model with cross-entropy loss alone. In setting 1 - 4, ‘Fine-tuning’ column refers to fine-tuning for $2$ to $5$ epochs after extracting subnetwork.
Table 5: Ablation study on CelebA (target: blond hair, spurious attribute: male). Effect of pruning (Pr) and finetuning (Ft) steps with (✓) and without (✗) our contrastive loss, cells marked ‘n.a.’ show no loss values because there is no corresponding pruning or fine-tuning stage. ✓last means fine-tuning last layer only.
Setting Fine-tuning ConLoss Pr ConLoss Ft WGA AVG UA
With Subnetwork Extraction
1 67.7 83.7 80.4
2 (PruSC) 89.6 90.1 90.5
3 41.8 91.5 73.9
4 n.a. 43.7 89.6 74.4
Without Subnetwork Extraction
5 n.a. 67.4 88.8 81.6
6 n.a. 78.9 87.9 87.2
7 (DFRclass)     ✓last n.a. 79.5 83.8 85.5
Subnetwork Extraction.
Settings $5$ to $7$ investigate whether one can improve the worst-group accuracy only by fine-tuning and take advantage of designing a class-balanced subset ($\mathcal{D}_{\text{task}}$) without a subnetwork. We skip the Binary Mask Training step (cf. Sec. 5.3) and either fine-tune the whole network (settings $5$ and $6$), or only the last layer (setting $7$) for the same number of epochs with $\mathcal{D}_{\text{task}}$. The latter setting is equivalent to DFRclass, with $\mathcal{D}_{\text{task}}$. We observe no significant improvement when fine-tuning the whole network compared to fine-tuning only the last layer (setting 6 vs. 7), fine-tuning the whole network even results in worse WGA.